# i-PI 3.0: a flexible, efficient framework for advanced atomistic simulations

### v1

@@ IV.3 Ensemble models and error estimation
The first entails independent models that are run as separate clients, by specifying multiple socket connectors as part of a <ffcommittee> tag. For instance,
<ffcommittee name="my_committee">
<ffsocket name=’driver-1’>
</ffsocket>
<ffsocket name=’driver-2’>
</ffsocket>
</ffcommittee>
The associated class takes care of collecting potentials, forces and stresses from the various sockets and averaging them. Even though a ffcommittee section contains multiple sockets, one should use a single <force> block referring to the name of the committee (my_committee in this example) in the definition of the system. A linear committee calibration scaling [136] can also be applied with the keyword <alpha>.
In the second mode, which is suitable when using a shallow ensemble model with weight sharing, a single client computes all the ensemble members. In this case, the client passes the average energies, forces and virials to i-PI  in the standard way and also passes the ensemble information as a JSON string into the extras field. The string contains the list of potentials, forces, and virials into fields named committee_pot, committee_force, committee_virial. We provide a reference implementation of such a client in the FORTRAN driver distributed with i-PI. It is then sufficient to specify a single client in the committee section
<ffcommittee name="my_committee">
<ffsocket name=’driver’>
</ffsocket>
<parse_json> True </parse_json>
</ffcommittee>
and to activate the flag parse_json that instructs the committee class to parse the ensemble data out of the string. Also in this case, the forces section should refer to "my_committee" and not to the name of the inner forcefield.
@@
Figure 12: Trajectories of the various energy predictions of the DPOSE committee model and of the order parameter that determines the pinning potential (11). The inset shows the near-Gaussian correlation plot between the order parameter and the deviation of one of the ensemble members from the mean potential.
@@ IV.3.1 Computing the solid-liquid chemical potential for water
We discuss a representative example of the use of ffcommittee to compute errors in on the solid-liquid chemical potential $\Delta\mu$ for an ice/water system, that showcases several of the advanced features in i-PI. This type of simulation can be used to estimate the melting point of water $T_{\text{m}}$, by repeating the procedure at different temperatures and finding the temperature at which $\Delta\mu=0$. As shown in Ref. 142, once a collection of $\Delta\mu$ values associated with the different ensemble members has been obtained, it is easy to determine the error on $T_{\text{m}}$ by ensemble propagation.
For this example, we use a Behler-Parrinello-style neural network,[143] based on smooth-overlap of atomic positions (SOAP)[144] features and trained on the dataset of Ref. 14. The model generates an ensemble of 5 energy predictions, differing only by the last-layer weights and trained by optimizing the negative log likelihood following the direct propagation of shallow ensembles (DPOSE) protocol.[100] We run a simulation with 336 water molecules, arranged in an elongated simulation cell containing coexisting water and ice. A restraining pinning potential to an anchor point, $\bar{n}$,
@@
is applied using a suitable collective variable that estimates the number of molecules in a solid-like environment, $n_{\text{S}}(A)$,[145] as implemented in PLUMED.[146] At the melting point, the potential restraints the solid fraction to the target value $\bar{n}$, whereas away from equilibrium the mean deviation is proportional to the solid-liquid molar free-energy difference, $\Delta\mu\propto\left<n_{\text{S}}(A)-\bar{n}\right>$. Further details on the setup can be inferred from the example files included in the demos/ensemble-deltamu folder. It should be noted that both size and duration of the simulations are insufficient to obtain a converged estimate of the melting point, as this is only intended as a demonstrative example.
@@
Figure 13: Cumulative sums of the reweighted mean $\langle n_{\text{S}}-\bar{n}\rangle_{k}$ computed using the $k$-th ensemble member. The dashed black line indicates the cumulative average of $n_{\text{S}}$ computed without reweighting. Panel a) shows the average computed with Boltzmann reweighting while panel b) uses a statistically-better-behaved cumulant expansion approximation (CEA).
@@
Fig. 12 shows a small segment of a trajectory obtained with the DPOSE model. The mean potential undergoes thermal fluctuations, and the ensemble of predictions has a varying spread along the trajectory. Similar to what was observed in Ref. 100, the members of the ensemble differ mostly by a rigid shift, reflecting a systematic bias in the prediction of the potential. Meanwhile, the order parameter fluctuates about a value that is slightly below the target, indicating that the simulation temperature (290 K) is above the melting point for this water model – at least for this under-converged supercell size. There is no obvious pattern in the correlation between the total potential energy and the order parameter: at this temperature, and for such a comparatively small system, random fluctuations dominate over the enthalpy change associated with the changes in the solid fraction. Importantly, the cross-correlations between the solid fraction and the ensemble potential shift $(V^{(k)}(A)-\bar{V}(A))$ show approximately Gaussian behaviour (see inset in Fig. 12). This simplifies the estimation of the average deviation $(n_{\text{S}}(A)-\bar{n})\right>$ from a single trajectory driven by $\bar{V}$. The basic idea is to use Torrie-Valleau reweighting,[147] i.e. weighting each snapshot by $e^{-\beta(V^{(k)}(A)-\bar{V}(A))}$. As shown in Fig. 13, the cumulative averages computed with the different weighting factors converge to different values – providing an ensemble of predictions for $\langle n_{\text{s}}-\bar{n}\rangle_{k}$. The sharp, sudden changes in the cumulative average reflect the occurrence of very large weighting factors, that occur whenever the deviations of individual predictions from the mean potential are large relative to $k_{\text{B}}T$. The resulting statistical inefficiency[87] would become worse as the system size is increased, eventually making the procedure completely ineffective. Following Ref. 142, one can use instead a cumulant expansion approximation (CEA) that estimates reweighted average under the assumption that the target observable and the potential are joint Gaussians, and yields
@@
This approximation introduces a systematic error because the distribution is not exactly Gaussian, but as can be seen in Fig. 12, the qualitative behaviour reflects the one seen for the direct reweighting. The statistical error is well-behaved also for large simulations, making the CEA a more generally applicable strategy. We provide a commented example to perform this kind of simulation in the demos/ensemble-deltamu; the post-processing needed to evaluate the reweighted averages is implemented in the i-pi-committee-reweight script. By repeating the simulation at different temperatures one can determine the temperature dependence of $\Delta\mu$ for each ensemble member separately, and use the separately-computed values of $T_{m}$ to estimate the error in the determination of the melting point. It is worth noting that in Ref. 142 the melting point was determined to be $290\pm 5$ K, with a setup similar to the one discussed here, but using a deep ensemble of models, fitted on the same dataset. This estimate is consistent with our observation that $\langle n_{\text{s}}-\bar{n}\rangle$ at 290 K is very close to zero. The discrepancy with Ref. 14, where the classical melting point for the reference DFT was determined by free-energy perturbation to be at $\approx 275$ K can be attributed to finite-size effects, that tend to de-stabilize the liquid and increase the melting point.
@@ IV.4 Cavity Molecular Dynamics for Polaritonics
For molecules confined in optical cavities, molecular transitions may form polaritons, the hybrid light-matter states, with cavity photon modes.[148, 149, 150, 151, 152] Of particular interest is the vibrational strong coupling regime,[153, 154] in which vibrational polaritons are formed when a molecular vibrational mode is near resonant with an IR cavity photon mode. Under vibrational strong coupling, many molecular properties, including chemical reaction rates and energy transfer rates, can be significantly modified in the liquid phase.[155, 156, 157] For a better understanding of these exciting experiments, both the molecules and the cavity photons need to be propagated self-consistently with realistic parameters. This requirement is beyond the scope of standard molecular dynamics schemes, whereby the time-dependent dynamics of the photon field are not explicitly propagated.
The i-PI package provides an efficient method for simulating vibrational strong coupling: the CavMD approach.[92, 93] Within the framework of CavMD, both nuclei and cavity photons in the infrared domain are treated classically, propagated by the following light-matter Hamiltonian under the ground electronic state: $H_{\rm QED}=H_{\rm M}+H_{\rm F}$,[158, 92] where $H_{\rm M}$ denotes the standard molecular Hamiltonian composed of the kinetic and potential terms, and
@@
The basic implementation strategy of CavMD is as follows. The cavity photon modes are stored in i-PI as L (light) atoms. The coordinates and velocities of both the nuclei and the cavity photon modes are propagated in i-PI. A new FFCavPhSocket class, which inherits from the original FFSocket class, evaluates the forces of both the nuclei and the cavity photon modes. Similar to FFSocket, FFCavPhSocket calls external client codes to evaluate the nuclear forces associated with $H_{\rm M}$, via the socket interface. After obtaining this component of the total forces, FFCavPhSocket performs additional calculations to obtain the photonic forces as prescribed by Eq. 13. Because this implementation encapsulates the cavity-related calculations within the FFCavPhSocket class, many advanced features in i-PI, including RPMD, [92] and ring-polymer instantons, can be directly applied to study vibrational strong coupling without additional modifications of the i-PI code.
Figure 14: CavMD simulations: Panel (a) illustrates the simulation setup of CavMD, whereby a large ensemble of molecules is confined in an optical cavity under vibrational strong coupling. The rest of the panels are adapted from Ref. 92 and show the IR spectra of the OH stretch mode of liquid water outside the cavity and inside the cavity with different light-matter coupling strength, $\widetilde{\varepsilon}$. LP and UP stand for lower and upper polariton, respectively.
The CavMD implementation was used to simulate the IR spectra of liquid water under vibrational strong coupling.[92] The key i-PI input section to simulate vibrational strong coupling is as follows:
<ffcavphsocket name=’lammps mode=’unix’>
<address>h2o-cl-cavmd</address>
<apply_photon> True </apply_photon>
<E0> 4e-4 </E0>
<omega_c units=’inversecm’> 3550.0 </omega_c>
<charge_array> [...] </charge_array>
</ffcavphsocket>
In the FFCavPhSocket section above, a cavity mode with frequency $\omega_{\rm c}=3550$ cm-1 was coupled to liquid water with an effective light-matter coupling strength $\widetilde{\varepsilon}=4\times 10^{-4}$ a.u. Under the default setting, the cavity mode is polarised along both the $x$ and $y$ directions. The partial charges of all molecular atoms are given in the charge_array section of FFCavPhSocket, and are used to evaluate the molecular dipole moment in Eq. (13). In Fig. 14 we show a schematic picture of the setup along with the IR spectra of water inside and outside the cavity. As shown in Fig. 14c, via classical CavMD simulations, the wide OH band of liquid water splits into a pair of polariton states, the upper polariton (UP) and the lower polariton (LP). Beyond this study, CavMD was also used to explore novel energy transfer and chemical reaction pathways under vibrational strong coupling. [159, 160, 161]
@@ V.1 Path Integral Coarse-Grained Simulations
Figure 15: IR, isotropic Raman, and anisotropic Raman vibrational spectrum of water at 300 K using the Te PIGS method. Adapted from Ref. 162.
@@
The imaginary-time path integral approach provides a rigorous framework for equilibrium time-independent properties but it doesn’t give access to real-time quantum dynamics.[163] Nevertheless, it is possible to extract meaningful dynamical information from path integral simulations by making approximations to Matsubara dynamics.[164] Standard path-integral based methods for approximate quantum dynamics, (thermostatted) ring polymer molecular dynamics [74, 78] (RPMD) and centroid molecular dynamics [77] (CMD), can be shown to be approximations to Matsubara dynamics.[165] Unfortunately, both RPMD and CMD exhibit accuracy and cost limitations. For instance, in estimating the vibrational density of states, RPMD, CMD and variants [166, 167, 168, 169] present artefacts such as unphysical shifted or broadened peaks.[170, 166] In addition, the computational overhead and impact of the artefacts become more pronounced with lowering temperature.
In their recent work,  Musil et al. [79] introduce a machine learning framework to significantly enhance the accuracy and efficiency of approximating quantum dynamics based on path-integral simulations. This approach is closely related to CMD, which involves the classical time-evolution of a system on the centroid potential of mean force, effectively capturing nuclear quantum effects up to the fluctuations of the path integral centroid. This framework has two components: a temperature elevation (Te) ansatz and path integral coarse-graining simulations (PIGS). The Te ansatz addresses issues such as spurious redshift in CMD, resulting in notable improvements in the accuracy of vibrational spectra in bulk [162] and interfacial [171] systems compared to existing methodologies. Meanwhile, PIGS enables the machine learning of the centroid potential of mean force using state-of-the-art machine learning potentials. Thus, the Te PIGS approach allows accurate modelling of quantum dynamics in condensed phase systems via simple and inexpensive classical MD on a machine learning potential.
This approach is easily implemented using the flexible functionality of i-PI as demonstrated for various vibrational spectroscopic observables of liquid water (see Fig. 15) using the MACE-OFF23(S) general organic machine learning potential.[162] For a given potential energy surface $V(\mathbf{q})$, training the Te PIGS machine learning potential requires data from an elevated-temperature simulation, such as the centroid positions $\bar{\mathbf{q}}=\frac{1}{P}~{}\sum_{j=1}^{P}\mathbf{q^{(j)}}$, the mean force acting on the centroid, $\mathbf{f}^{\text{CMD}}=-\frac{1}{P}~{}\sum_{j=1}^{P}\frac{\partial V(\mathbf{q}^{(j)})}{\partial\mathbf{q}^{(j)}}$, and the physical force acting on the centroid, $\mathbf{f}=-\frac{\partial V(\bar{\mathbf{q})}}{\partial\bar{\mathbf{q}}}$. In this example from Ref. 162, the elevated temperature is set to $T_{e}$ = 500 K and the centroid potential of mean force is fit from a 10 ps long path integral simulation requiring just 8 replicas, with the centroid positions and forces sampled every 10 fs (20 timesteps). The centroid positions and forces can be printed out as follows:
<trajectory filename=’xc format=’xyz
stride=’20’> x_centroid </trajectory>
<trajectory filename=’fc format=’xyz
stride=’20’> f_centroid </trajectory>
To calculate the physical force acting on the centroid positions, we perform a “replay” simulation, which allows the calculation and output of properties or trajectory information on a list of centroid positions using:
<trajectory filename=’for format=’xyz
stride=’1’> forces </trajectory>
<motion mode=’replay’>
<file mode=’xyz’> simulation.xc.xyz </file>
</motion>
The Te PIGS machine learning potential can be developed using standard packages like n2p2 [172], nequip [173], mace [174] using so-called “delta learning” by regressing on the centroid positions as the input and the difference between $\mathbf{f}^{\text{CMD}}$ and $\mathbf{f}$ as the target.
Performing a Te PIGS simulation only requires classical molecular dynamics on the sum of the potential energy surface $U(\mathbf{q})$ and the Te PIGS machine learning potential. This can done in i-PI by exploiting the modularity of forces. As shown in the following snippet, this setup requires to define two forcefield sockets for the potential energy surface and the Te PIGS machine learning potential:
<ffsocket name=’PES mode=’unix’>
<address> driver.1 </address>
</ffsocket>
<ffsocket name=’TEPIGS mode=’unix’>
<address> driver.2 </address>
</ffsocket>
The addition of the potentials along with the calculation of all derivatives like the force and stress is done under the hood by defining the forces of the system class comprising two force components that communicate with the forcefield sockets:
<forces>
<force forcefield=’PES weight=’1.0’>
</force>
<force forcefield=’TEPIGS weight=’1.0’>
</force>
</forces>
@@
Nonlinear, multi-time response functions are associated with different flavours of time-resolved and multidimensional spectroscopic techniques.[175] Here, we show an example provided in i-PI demos, that simulates the two-dimensional IR-Raman spectrum of liquid water in which the sample interacts first with two mid- or far-IR pulses through a dipole interaction and then with a near-IR or visible pulse through a Raman process.[176, 177] The spectrum is obtained as a two-dimensional Fourier sine transform of the two-time response function
@@
where $\hat{\alpha}$ is the polarizability operator and $\hat{\mu}$ the dipole moment. The example presented in demos/2D-IR-Raman implements the equilibrium-nonequilibrium RPMD method of Refs. 95-96, which approximates the true quantum-mechanical response function (14). As a special case when the number of ring-polymer beads $P$ is 1, the method corresponds to the MD approach of Hasegawa and Tanimura.[94] In passing, we note that other methods based on imaginary-time path integrals have been reported to calculate non-linear response functions using approximations to fully-symmetrized multi-time Kubo correlation functions.[178, 179, 180]
@@
Figure 16: Two-dimensional IR-Raman spectroscopy: Panel (a) shows the schematic of the equilibrium-nonequilibrium (RP)MD approach (see text). Panel (b) and (c) depict the frequency-dependent response function corresponding to the two-dimensional THz-IR-Raman (also called THz-IR-visible [176]) spectrum of liquid water, computed using classical MD and TRPMD, respectively. Figure adapted from Ref. 96.

### v2

@@ IV.3 Ensemble models and error estimation
The first entails independent models that are run as separate clients, by specifying multiple socket connectors as part of a <ffcommittee> tag. For instance, <ffcommittee name=’my_committee’> <ffsocket name=’driver-1’> [...] </ffsocket> <ffsocket name=’driver-2’> [...] </ffsocket> [...] </ffcommittee> The associated class takes care of collecting potentials, forces and stresses from the various sockets and averaging them. Even though a ffcommittee section contains multiple sockets, one should use a single <force> block referring to the name of the committee (my_committee in this example) in the definition of the system. A linear committee calibration scaling [162] can also be applied with the keyword <alpha>.
In the second mode, which is suitable when using a shallow ensemble model with weight sharing, a single client computes all the ensemble members. In this case, the client passes the average energies, forces and virials to i-PI  in the standard way and also passes the ensemble information as a JSON string into the extras field. The string contains the list of potentials, forces, and virials into fields named committee_pot, committee_force, committee_virial. We provide a reference implementation of such a client in the FORTRAN driver distributed with i-PI. It is then sufficient to specify a single client in the committee section <ffcommittee name=’my_committee’> <ffsocket name=’driver’> [...] </ffsocket> <parse_json> True </parse_json> </ffcommittee> and to activate the flag parse_json that instructs the committee class to parse the ensemble data out of the string. Also in this case, the forces section should refer to "my_committee" and not to the name of the inner forcefield.
@@
Figure 12: Trajectories of the various energy predictions of the DPOSE [115] committee model (plotted in constrasting colors) and of the order parameter that determines the pinning potential, Eq. 11 (black). The inset shows the near-Gaussian correlation plot between the order parameter and the deviation of one of the ensemble members from the mean potential.
@@ IV.3.1 Computing the solid-liquid chemical potential for water
We discuss a representative example of the use of ffcommittee to compute errors of the solid-liquid chemical potential $\Delta\mu$ for an ice/water system, that showcases several of the advanced features in i-PI. This type of simulation can be used to estimate the melting point of water $T_{\text{m}}$, by repeating the procedure at different temperatures and finding the temperature at which $\Delta\mu=0$. As shown in Ref. 107, once a collection of $\Delta\mu$ values associated with the different ensemble members has been obtained, it is easy to determine the error on $T_{\text{m}}$ by ensemble propagation.
For this example, we use a Behler-Parrinello-style neural network,[168] based on smooth-overlap of atomic positions (SOAP)[169] features and trained on the dataset of Ref. 14. The model generates an ensemble of 5 energy predictions, differing only by the last-layer weights and trained by optimizing the negative log-likelihood following the direct propagation of shallow ensembles (DPOSE) protocol.[115] We run a simulation with 336 water molecules, arranged in an elongated simulation cell containing coexisting water and ice. A restraining pinning potential to an anchor point, $\bar{n}$,
@@
is applied using a suitable collective variable that estimates the number of molecules in a solid-like environment, $n_{\text{S}}(A)$,[170] as implemented in PLUMED[39] and here used through its interface with i-PI. At the melting point, the potential restraints the solid fraction to the target value $\bar{n}$, whereas away from equilibrium the mean deviation is proportional to the solid-liquid molar free-energy difference, $\Delta\mu\propto\left<n_{\text{S}}(A)-\bar{n}\right>$. Further details on the setup can be inferred from the example files included in the demos/ensemble-deltamu directory. It should be noted that both size and duration of the simulations are insufficient to obtain a converged estimate of the melting point, as this is only intended as a demonstrative example.
@@
Figure 13: Cumulative sums of the reweighted mean $\langle n_{\text{S}}-\bar{n}\rangle_{k}$ computed using the $k$-th ensemble member (plotted in contrasting colors). The dashed black line indicates the cumulative average of $n_{\text{S}}$ computed without reweighting. Panel a) shows the average computed with Boltzmann reweighting while panel b) uses a statistically-better-behaved cumulant expansion approximation (CEA).
@@
Fig. 12 shows a small segment of a trajectory obtained with the DPOSE model[115]. The mean potential undergoes thermal fluctuations, and the ensemble of predictions has a varying spread along the trajectory. Similar to what was observed in Ref. 115, the members of the ensemble differ mostly by a rigid shift, reflecting a systematic bias in the prediction of the potential. Meanwhile, the order parameter fluctuates about a value that is slightly below the target, indicating that the simulation temperature (290 K) is above the melting point for this water model – at least for this under-converged supercell size. There is no obvious pattern in the correlation between the total potential energy and the order parameter: at this temperature, and for such a comparatively small system, random fluctuations dominate over the enthalpy change associated with the changes in the solid fraction. Importantly, the cross-correlations between the solid fraction and the ensemble potential shift $(V^{(k)}(A)-\bar{V}(A))$ show approximately Gaussian behaviour (see inset in Fig. 12). This simplifies the estimation of the average deviation $(n_{\text{S}}(A)-\bar{n})\right>$ from a single trajectory driven by $\bar{V}$. The basic idea is to use Torrie-Valleau reweighting,[171] i.e. weighting each snapshot by $e^{-\beta(V^{(k)}(A)-\bar{V}(A))}$.
As shown in Fig. 13, the cumulative averages computed with the different weighting factors converge to different values – providing an ensemble of predictions for $\langle n_{\text{s}}-\bar{n}\rangle_{k}$. The sharp, sudden changes in the cumulative average reflect the presence of very large weighting factors, that occur whenever the deviations of individual predictions from the mean potential are large relative to $k_{\text{B}}T$. The resulting statistical inefficiency[101] would become worse as the system size is increased, eventually making the procedure completely ineffective. Following Ref. 107, one can use instead a cumulant expansion approximation (CEA) that estimates reweighted average under the assumption that the target observable and the potential are joint Gaussians, and yields
@@
This approximation introduces a systematic error because the distribution is not exactly Gaussian, but as can be seen in Fig. 12, the qualitative behaviour reflects the one seen for the direct reweighting. The statistical error is well-behaved also for large simulations, making the CEA a more generally applicable strategy. We provide a commented example to perform this kind of simulation in the demos/ensemble-deltamu; the post-processing needed to evaluate the reweighted averages is implemented in the i-pi-committee-reweight script.
By repeating the simulation at different temperatures one can determine the temperature dependence of $\Delta\mu$ for each ensemble member separately, and use the separately-computed values of $T_{m}$ to estimate the error in the determination of the melting point. It is worth noting that in Ref. 107 the melting point was determined to be $290\pm 5$ K, with a setup similar to the one discussed here, but using a deep ensemble of models, fitted on the same dataset. This estimate is consistent with our observation that $\langle n_{\text{s}}-\bar{n}\rangle$ at 290 K is very close to zero. The discrepancy with Ref. 14, where the classical melting point for the reference DFT was determined by free-energy perturbation to be at $\approx 275$ K can be attributed to finite-size effects, that tend to destabilise the liquid and increase the melting point.
@@ IV.4 Cavity Molecular Dynamics for Polaritonics
For molecules confined in optical cavities, molecular transitions may form polaritons, hybrid light-matter states with cavity photon modes.[172, 173, 174, 175, 176] Of particular interest is the vibrational strong coupling regime,[177, 178] in which vibrational polaritons are formed when a molecular vibrational mode is near resonant with an IR cavity photon mode. Under vibrational strong coupling, many molecular properties, including chemical reaction rates and energy transfer rates, can be significantly modified in the liquid phase.[179, 180, 181] For a better understanding of these exciting experiments, both the molecules and the cavity photons need to be propagated self-consistently with realistic parameters. This requirement is beyond the scope of standard molecular dynamics schemes, whereby the time-dependent dynamics of the photon field are not explicitly propagated.
The 3.0 version of the i-PI package provides an efficient method for simulating vibrational strong coupling: the CavMD approach.[111, 112] Within the framework of CavMD, both nuclei and cavity photons in the IR domain are treated classically, propagated by the following light-matter Hamiltonian under the ground electronic state: $H_{\rm QED}=H_{\rm M}+H_{\rm F}$,[182, 111] where $H_{\rm M}$ denotes the standard molecular Hamiltonian composed of the kinetic and potential terms, and
@@
Figure 14: CavMD simulations: Panel (a) illustrates the simulation setup of CavMD, whereby a large ensemble of molecules is confined in an optical cavity under vibrational strong coupling. The rest of the panels are adapted from Ref. 111, with permission from the authors, and show the IR spectra of the OH stretch mode of liquid water without a cavity and inside a cavity with different light-matter coupling strength, $\widetilde{\varepsilon}$. LP and UP stand for lower and upper polariton, respectively.
The basic implementation strategy of CavMD is as follows. The cavity photon modes are stored in i-PI as L (light) atoms. The coordinates and velocities of both the nuclei and the cavity photon modes are propagated in i-PI. A new FFCavPhSocket class, which inherits from the original FFSocket class, evaluates the forces of both the nuclei and the cavity photon modes. Similar to FFSocket, FFCavPhSocket calls external client codes to evaluate the nuclear forces associated with $H_{\rm M}$, via the socket interface. After obtaining this component of the total forces, FFCavPhSocket performs additional calculations to obtain the photonic forces and cavity modifications on the nuclear forces as prescribed by Eq. 13. Because this implementation encapsulates the cavity-related calculations within the FFCavPhSocket class, many advanced features in i-PI, including RPMD, [111] and ring-polymer instantons, can be directly applied to study vibrational strong coupling without additional modifications of the i-PI code.
The CavMD implementation was used to simulate the IR spectra of liquid water under vibrational strong coupling.[111] The key i-PI input section to simulate vibrational strong coupling is as follows: <ffcavphsocket name=’lammps mode=’unix’> <address>h2o-cl-cavmd</address> <apply_photon> True </apply_photon> <E0> 4e-4 </E0> <omega_c units=’inversecm’> 3550.0 </omega_c> <charge_array> [...] </charge_array> </ffcavphsocket> In the FFCavPhSocket section above, a cavity mode with frequency $\omega_{\rm c}=3550$ cm-1 was coupled to liquid water with an effective light-matter coupling strength $\widetilde{\varepsilon}=4\times 10^{-4}$ a.u. Under the default setting, the cavity mode is polarised along both the $x$ and $y$ directions. The partial charges of all molecular atoms are given in the charge_array section of FFCavPhSocket, and are used to evaluate the molecular dipole moment in Eq. (13). In Fig. 14 we show a schematic picture of the setup along with the IR spectra of water inside and outside the cavity. As shown in Fig. 14c, via classical CavMD simulations, the wide OH band of liquid water splits into a pair of polariton states, the upper polariton (UP) and the lower polariton (LP). Beyond this study, CavMD was also used to explore novel energy transfer and chemical reaction pathways under vibrational strong coupling. [183, 184, 185]
@@ V.1 Path Integral Coarse-Grained Simulations
[…]
@@
While PIMD has demonstrated success in incorporating quantum nuclear motion for equilibrium properties, extracting accurate dynamical information from path-integral simulations is a longstanding challenge [186]. For instance, in estimating the vibrational spectra, established path-integral approximations to quantum dynamics such as with RPMD [90], CMD [92, 93] and variants [94, 75, 187, 188, 189], present artefacts such as unphysical shifted or broadened peaks.[190, 94] In addition, the computational overhead of path-integral simulations and the impact of the artefacts become more pronounced with lowering temperature.
[…]
The quantity $\bar{\mathbf{f}}_{\text{p}}$ can be calculated in a post-processing step with a “replay” simulation, which allows the calculation and output of properties on a list of centroid positions. Alternatively, the physical force on the centroid can be printed on the fly by defining a “dummy” force component with a zero weight that receives centroid positions using, <forces> [...] <force forcefield=’maceoff23 weight=’0’ nbeads=’1’> </force> </forces> Printing out this force component requires a single line: <trajectory filename=’physical_force stride=’20’ format=’ase’> forces_component_raw(1) </trajectory>
[…]
Figure 16: Two-dimensional IR-Raman spectroscopy: Panel (a) shows the schematic of the equilibrium-nonequilibrium (RP)MD approach (see text). Panel (b) and (c) depict the frequency-dependent response function corresponding to the two-dimensional THz-IR-Raman (also called THz-IR-visible [194]) spectrum of liquid water, computed using classical MD and TRPMD, respectively. Figure adapted from Ref. 119, with permission (CC-BY Licence).
@@
Nonlinear, multi-time response functions are associated with different flavours of time-resolved and multidimensional spectroscopic techniques.[195] Here, we show an example provided in the i-PI demos directory, that simulates the two-dimensional IR-Raman spectrum of liquid water in which the sample interacts first with two mid- or far-IR pulses through a dipole interaction and then with a near-IR or visible pulse through a Raman process.[194, 196] The spectrum is obtained as a two-dimensional Fourier sine transform of the two-time response function
@@
where $\hat{\alpha}$ is the polarizability operator and $\hat{\mu}$ the dipole moment. The example presented in demos/2D-IR-Raman implements the equilibrium-nonequilibrium RPMD method of Refs. 118 and 119, which approximates the quantum-mechanical response function of Eq. 14. As a special case when the number of ring-polymer beads $P=1$, the method corresponds to the MD approach of Hasegawa and Tanimura.[117] In passing, we note that other methods based on imaginary-time path integrals have been reported to calculate non-linear response functions using approximations to fully-symmetrized multi-time Kubo correlation functions.[197, 198, 199]
@@
Figure 17: (a) Ratio of the structure factors obtained from classical ($S_{\mathrm{CL}}$) and quantum path-integral ($S_{\mathrm{PI}}$) simulations $S_{\text{CL}}/S_{\text{PI}}$ performed on a $18\times 18$ supercell of monolayer 1H-TaS2. A value close to 1 (indicated in white) corresponds to minimal differences between classical and quantum simulations. The peaks at $\bm{q}=2/3\Gamma M$ and $\bm{q}=K$ for $T=50$ K are characteristic of the $3\times 3$ CDW phase. (b) Temperature-dependent structure factor $\langle S(\bm{q})\rangle_{T}$ at the characteristic CDW wavevector $\bm{q}=2/3\Gamma M$ obtained from classical MD replica-exchange MD (light blue) and path integral replica-exchange MD (dark blue). The experimental value (for the bulk) is represented by the vertical black dashed line. The shift of the inflection point of the PIMD curve toward the experimental value is due to nuclear quantum effects. Figure and caption adapted from Ref. 20 (CC-BY Licence).